### v1

# Lookahead-then-Verify: Reliable Constrained Decoding for Diffusion LLMs under Context-Free Grammars

Abstract.
Diffusion Large Language Models (dLLMs) have demonstrated promising generative capabilities and are increasingly used to produce formal languages defined by context-free grammars, such as source code and chemical expressions. However, as probabilistic models, they still struggle to generate syntactically valid outputs reliably. A natural and promising direction to address this issue is to adapt constrained decoding techniques to enforce grammatical correctness during generation. However, applying these techniques faces two primary obstacles. On the one hand, the non-autoregressive nature of dLLMs renders most existing constrained decoding approaches inapplicable. On the other hand, current approaches specifically designed for dLLMs may allow intermediate outputs that are impossible to complete into valid sentences, which significantly limits their reliability in practice.
To address these challenges, we present LAVE, a constrained decoding approach specifically designed for dLLMs. Our approach leverages a key property of dLLMs, namely their ability to predict token distributions for all positions in parallel during each forward pass. Whenever a new token is proposed by model, LAVE performs lookahead using these distributions to efficiently and reliably verify the validity of the proposed token. This design ensures reliable constraints by reliably preserving the potential for intermediate outputs to be extended into valid sentences. Extensive experiments across four widely used dLLMs and three representative benchmarks demonstrate that LAVE consistently outperforms existing baselines and achieves substantial improvements in syntactic correctness, while incurring negligible runtime overhead.
1. Introduction
Recently, diffusion Large Language Models (dLLMs) have attracted significant attention in both academia and industry (Li et al., 2025a, c; Cheng et al., 2025; Liu et al., 2025; Song et al., 2025). Unlike AutoRegressive LLMs (AR LLMs), dLLMs generate tokens in a non-autoregressive manner, which allows tokens on the right to be generated before those on the left, as shown in Figure 1. Due to their promising capabilities, they are increasingly applied to tasks requiring the generation of formal languages, which are typically defined by a Context-Free Grammar (CFG) (Sun et al., 2025; Ugare et al., 2024c). For instance, a growing body of research has explored the potential of dLLMs in generating source code (Gong et al., 2025; Xie et al., 2025) and chemical expressions (Mündler et al., 2025). Moreover, the industry has begun adopting dLLMs, such as Gemini Diffusion (DeepMind, 2025) and Mercury Coder (Khanna et al., 2025), for tasks including code generation. However, as probabilistic models, dLLMs still struggle to generate syntactically valid outputs reliably in formal languages (Yang et al., 2025a; Suresh et al., 2025), which limits their broader applicability in real-world scenarios. To assess the severity of this issue, we conducted a preliminary experiment on HumanEval-CPP (Zheng et al., 2023), a widely used benchmark for evaluating the code generation capabilities of LLMs. Our results show that even leading dLLMs suffer from high syntax error rates. For example, Dream-v0-Instruct-7B (Ye et al., 2025) exhibits a syntax error rate of up to 23.8%.
To address this reliability issue, a promising direction is to adapt constrained decoding techniques, which have been widely used to ensure that AR LLMs produce syntactically valid outputs (Willard and Louf, 2023; Chen et al., 2025; Beurer-Kellner et al., 2024; Ugare et al., 2024c; Wang et al., 2024). These approaches typically rely on grammar parsers (Aho and Johnson, 1974; DeRemer, 1971; Earley, 1970) to guarantee the validity of each newly generated token (Wagner and Graham, 1998; Microsoft, 2025; Willard and Louf, 2023). Existing parsers are generally restricted to processing complete prefixes (i.e., prefixes containing no ungenerated token), which naturally aligns with the autoregressive generation process of AR LLMs (Sun et al., 2025; DeRemer, 1971; Geng et al., 2023). However, dLLMs generate tokens in a non-autoregressive manner. As shown in Figure 1(b), dLLMs often produce incomplete prefixes (i.e., prefixes containing some ungenerated tokens – [MASK]), which renders existing constrained decoding techniques inapplicable. Consequently, a key challenge lies in how to enable constrained decoding to function effectively on such incomplete prefixes. Although recent work (Mündler et al., 2025) has attempted to design CFG-based constrained decoding for dLLMs, it fails to ensure that the constraints are reliable. More concretely, it often leads to intermediate generated outputs that are impossible to complete into a valid sentence in the target language, which significantly undermines its reliability. Therefore, developing constrained decoding approaches tailored to dLLMs still remains an open and pressing problem.
(a) Autoregressive LLM
(b) Diffusion LLM
Figure 1. AR LLMs generate tokens autoregressively, so every intermediate output forms a complete prefix, whereas dLLMs generate tokens non-autoregressively, yielding intermediate outputs that are often incomplete prefixes due to remaining [MASK].
To address existing challenge, we propose LAVE, a constrained decoding approach specifically designed for dLLMs. At the core of our approach is a simple but powerful idea: lookahead-then-verify. Unlike AR LLMs, which predict a single probability distribution for the next token, dLLMs predict distributions for all positions in parallel during every forward pass (Gong et al., 2024; Gao et al., 2025; Yang et al., 2025a). Leveraging this capability, when the model proposes a token for a masked position, LAVE can naturally perform lookahead by sampling tokens for the remaining masked positions within the current incomplete prefix. This process effectively transforms the incomplete prefix into a set of complete prefixes that reflect the likely continuations of the model and can be directly validated by most parsers (Earley, 1970; Aho and Johnson, 1974; DeRemer, 1971). Subsequently, LAVE employs a grammar parser to verify whether any of these lookahead complete prefixes can be further extended into a valid sentence in the target language. The existence of such a prefix serves as a reliable guarantee that the proposed token can preserve the potential for the output to be completed into a valid sentence. If this condition is satisfied, the proposed token is accepted; otherwise, it is rejected and the model is required to propose another token.
However, implementing this idea presents two potential computational hurdles. ❶ First, an incomplete prefix may contain multiple masked positions, and the vocabulary size is typically very large. As a result, determining whether a proposed token should be accepted may in principle require exploring an exponential number of lookahead complete prefixes, which is computationally infeasible. Fortunately, our preliminary experiments in Section 3.2.2 show that sampling only a small number of lookahead prefixes is usually sufficient to decide whether a proposed token is acceptable, which reduces the runtime overhead of verification to a negligible level. ❷ Second, even with efficient verification, the iterative proposal and verification process may occasionally become trapped in difficult contexts, where the model repeatedly proposes tokens that fail verification. To address this issue, LAVE incorporates a lightweight recovery mechanism in Section 3.3 that actively adjusts the current context, allowing the model to escape such stalled states while preserving constraint reliability.
Our experimental results demonstrate that LAVE consistently outperforms all baselines across four widely used dLLMs and three representative benchmarks covering diverse formal languages. Across nearly all experimental settings, LAVE improves the syntactic correctness rate to almost 100%, substantially enhancing the reliability of the model. Furthermore, in many scenarios, it yields improvements in functional correctness that are multiples of the gains achieved by other approaches. In terms of cost, we find that the runtime overhead introduced by LAVE is nearly negligible. We further conduct ablation studies, which show that each component contributes positively to the overall performance of LAVE. In addition, LAVE exhibits strong robustness to variations in key hyperparameters and consistently improves the quality of generated outputs across diverse settings.
In summary, our contributions are threefold:
  • We identify an unreliability issue in existing approaches for constraining dLLMs under CFGs, where generated tokens may render the output impossible to complete into a valid sentence.
  • We propose LAVE, a constrained decoding approach that reliably enforces CFG constraints for dLLMs, ensuring that intermediate outputs remain syntactically extendable.
  • We conduct extensive experiments using four widely used dLLMs on three representative benchmarks spanning multiple domains. The results demonstrate that LAVE substantially improves the reliability of dLLMs in generating CFG-compliant outputs.
2. Background and Related Work
In this section, we provide the necessary background to understand our approach and discuss related work.
2.1. Diffusion Large Language Models
Diffusion LLMs have recently emerged as a promising alternative to traditional autoregressive LLMs, offering advantages in flexible and parallel token prediction (Li et al., 2025a; Sahoo et al., 2024; Li et al., 2025b). Recent advancements have significantly narrowed the performance gap between dLLMs and AR LLMs, paving the way for the practical application of dLLMs in real-world scenarios. Early open-source dLLMs, such as LLaDA (Nie et al., 2025) and Dream (Ye et al., 2025), have demonstrated competitive performance comparable to LLaMA3-8B (Touvron et al., 2023). Furthermore, specialized models like Dream-Coder (Xie et al., 2025) and DiffuCoder (Gong et al., 2025) have been introduced to excel in formal languages such as source code, achieving results on parity with Qwen2.5-Coder-7B (Hui et al., 2024) across various code generation benchmarks. Mercury Coder (Khanna et al., 2025) and Gemini Diffusion (DeepMind, 2025) have attained capabilities rivaling those of proprietary commercial AR LLMs while offering superior inference speeds. More recently, LLaDA2 (Bie et al., 2025) has successfully scaled diffusion language models to 100 billion parameters and incorporates advanced functionalities such as function calling.
The inference process of mainstream dLLMs follows an iterative denoising procedure in which a fully masked sequence $\mathbf{y}^{0}$ is gradually transformed into the final output. Formally, a dLLM $f$ introduces a special [MASK] token and initializes generation with every position masked, denoted as $\mathbf{y}^{0}=(\texttt{[MASK]})_{i=1}^{L}$, where $L$ is a predefined max generation length. Given a prompt $\mathbf{p}$, the model performs up to $T$ denoising steps to produce the final sequence $\mathbf{y}^{T}=(y_{i}^{T})_{i=1}^{L}$, which contains no remaining [MASK] tokens. In practice, the number of denoising steps $T$ is typically no greater than $L$. At each step $t\in\{1,\dots,T\}$, the model performs a single forward pass to predict the token probability distributions $P_{i}(\cdot\mid\mathbf{p}\oplus\mathbf{y}^{t-1})$ in parallel for all positions $i$ that are still filled with [MASK] tokens. A small subset of these masked positions is then decoded into non-[MASK] tokens according to the predicted distributions. Crucially, the generation order does not strictly follow a left-to-right order. For instance, as shown in Figure 1(b), masks at later positions may be transformed into non-[MASK] tokens prior to those at earlier positions.
2.2. Constrained Decoding
Constrained decoding (Willard and Louf, 2023; Earley, 1970; Microsoft, 2025; Beurer-Kellner et al., 2024; Chen et al., 2025; Melcer et al., 2024; Park et al., 2025; Ugare et al., 2024b; Nakshatri et al., 2025; Ugare et al., 2024c; Loula et al., 2025; Geng et al., 2023) is widely used to ensure that the output generated by LLMs adheres to a given grammar, typically a Context-Free Grammar (CFG). However, most existing approaches in this line of work are specifically designed for AR LLMs and rely heavily on their left-to-right generation order. As a result, these approaches are not directly applicable to dLLMs, whose generation paradigm is inherently non-autoregressive. More recently, several studies have begun to explore constrained decoding tailored to dLLMs. For example, DINGO (Suresh et al., 2025) introduces a dynamic programming-based decoding procedure that enables dLLMs to generate grammar-compliant outputs. However, this approach is limited to regular grammars, which significantly restricts its applicability in realistic scenarios that typically require more expressive context-free grammars (Microsoft, 2025; Earley, 1970).
The work most closely related to ours is that of Mündler et al. (Mündler et al., 2025), who propose the first constrained decoding approach that enforces CFG compliance during dLLM generation. Their approach adopts a classical rejection sampling strategy (Parys et al., 2025; Jiang et al., 2024) to impose grammatical constraints. The key idea is to reduce the validity check for a newly proposed token to a language intersection problem. To illustrate, consider a model generating C code with the current intermediate output for ( [MASK] [MASK] …. If the model proposes the token ) for the second masked position, their approach constructs a regular language that accepts all strings matching the pattern for ( $\Sigma^{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}1}$ ) $\Sigma^{*}$, where $\Sigma$ denotes the vocabulary. This regular language is then intersected with the target context-free language, and the proposed token is accepted if and only if this intersection is non-empty.
Definition: Reliability of Constraints In this paper, we formally define the reliability of constraints as follows  [^1] Following most prior work (Willard and Louf, 2023; Beurer-Kellner et al., 2024; Ugare et al., 2024c) on constrained decoding for AR LLMs, we define reliability without taking the maximum generation length into account.. Reliable Constraint: At every step of decoding, the current intermediate output remains extendable, meaning that it can still be completed into at least one valid sentence in the target language. Unreliable Constraint: There exists at least one decoding step at which the intermediate output is not extendable, meaning that it can no longer be completed into any valid sentence in the target language. Since constrained decoding aims to improve the reliability of LLM generated output, enforcing a reliable constraint is a fundamental requirement for practical deployment.
To ensure computational tractability, their approach has to construct the regular language by ignoring the actual length limit of masked spans, assuming that each masked span can be filled with an arbitrary number of tokens. In the example above, the regular language will be over-approximated as for ( $\Sigma^{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}*}$ ) $\Sigma^{*}$ in practice. We argue that this over-approximation introduces substantially constraint unreliability, as it may accept proposed tokens that render the intermediate output impossible to complete into a valid sequence. For instance, revisiting the example above, there is only a single masked position between the parentheses, while a syntactically valid for statement requires at least two semicolons, which typically correspond to two tokens. Despite this requirement, the over-approximation allows the approach to accept the proposed token ), even though the resulting intermediate output can never be completed into a valid program.
Summary. Due to the non-autoregressive generation nature of dLLMs, the vast majority of existing CFG-based constrained decoding approaches are rendered inapplicable. Moreover, the current approach tailored for dLLMs fails to enforce reliable constraints during inference, which substantially undermines the reliability. This highlights the necessity of proposing a reliable constrained decoding approach tailored to dLLMs.
3. Methodology
In this section, we propose LAVE, a reliable constrained decoding approach designed for dLLMs. We first provide an overview of the LAVE in Section 3.1, followed by detailed descriptions of its key components and implementation details in Section 3.2 and Section 3.3.
Algorithm 1 LAVE
1:dLLM model $f$, Context-Free Grammar $\mathcal{G}$, prompt $\mathbf{p}$, lookahead size $N$, attempt budget $\tau$
2:Initialize $\mathbf{y}$ with [MASK] tokens
3:Initialize counter $c_{\text{fail}}\leftarrow 0$
4:Initialize $\mathbf{y}_{\text{cache}}\leftarrow\epsilon$
5:while existing remaining [MASK] do
6:  Propose a token $t^{*}$ for a masked position $i^{*}$ using $f$
7:  Update temporary updated output $\mathbf{y}^{\prime}\leftarrow\mathbf{y}$ with $\mathbf{y}^{\prime}_{i^{*}}\leftarrow t^{*}$
8:  if Lookahead-Based Verification($\mathbf{y}^{\prime}$) == true then $\triangleright$ Section 3.2
9:    $\mathbf{y}\leftarrow\mathbf{y}^{\prime}$ $\triangleright$ Accept the proposed token
10:    Update $\mathbf{y}_{\text{cache}}$ with a verified complete prefix
11:    $c_{\text{fail}}\leftarrow 0$
12:  else
13:    $c_{\text{fail}}\leftarrow c_{\text{fail}}+1$ $\triangleright$ Reject the proposed token
14:  end if
15:  if $c_{\text{fail}}\geq\tau$ then $\triangleright$ Section 3.3
16:    Apply Cache-Enhanced Recovery to $\mathbf{y}$ with $\mathbf{y}_{\text{cache}}$
17:    $c_{\text{fail}}\leftarrow 0$
18:  end if
19:end while
20:Output: final output $y$ without [MASK] tokens
3.1. Overview of LAVE
We constrain the decoding process of dLLMs through an iterative proposal and verification mechanism, as illustrated in Algorithm 1. During inference, the dLLM is allowed to propose new tokens for any masked positions according to its native decoding strategy. Each proposed token updates the current output by replacing a [MASK] token with a non-[MASK] token. The effectiveness of this decoding process critically depends on how the validity of each proposed token is assessed. To this end, we apply Lookahead-Based Verification (Section 3.2), the core component of LAVE, to reliably determine whether a proposed token should be accepted or rejected. If the token is accepted, the dLLM proceeds to generate the next token; otherwise, it attempts to propose an alternative.
In certain cases, the model may repeatedly fail to produce an acceptable proposal (Mündler et al., 2025; Jiang et al., 2024; Ugare et al., 2024a; Melcer et al., 2024). To mitigate this issue, we track the number of consecutive failed proposal attempts and invoke Cache-Enhanced Recovery (Section 3.3) once this number exceeds a predefined threshold, enabling the model to recover from stalled inference states.
3.2. Lookahead-Based Verification
In this section, we introduce Lookahead-Based Verification, which is the core of LAVE. Figure 2 illustrates how verification is performed.
When the model proposes a new token, it updates the current output by replacing a specific [MASK] token with a non-[MASK] token. The goal of Lookahead-Based Verification is to reliably verify whether this update preserves the potential for the output to be completed into a valid sentence.
3.2.1. Problem Formulation.
Formally, let $\mathbf{y}\in(\Sigma\cup\{\texttt{[MASK]}\})^{L}$ denote the updated output that includes the newly proposed token. We define the index of the rightmost non-[MASK] token as:
(1) $r(\mathbf{y})=\max\{\,i\mid y_{i}\neq\texttt{[MASK]}\,\},$
where $r(\mathbf{y})=0$ if all tokens are masked. This index determines a prefix $\mathbf{y}^{p}=\mathbf{y}_{1:r(\mathbf{y})}$, which is typically an incomplete prefix containing several [MASK] tokens. We denote the set of masked positions inside this prefix as:
(2) $\mathcal{M}(\mathbf{y}^{p})=\{\,i\mid 1\leq i<r(\mathbf{y}),\ y^{p}_{i}=\texttt{[MASK]}\,\}.$
A proposed token can be reliably accepted as long as there exists at least one assignment of tokens to these remaining masked positions such that the resulting complete prefix is a valid prefix:
(3) $\exists\mathbf{z}\in\Sigma^{|\mathcal{M}(\mathbf{y}^{p})|}\quad\text{s.t.}\quad\mathsf{IsExtendable}\!\left(\Phi(\mathbf{y}^{p},\mathbf{z})\right)=\texttt{true},$
where $\Phi(\cdot,\cdot)$ is a function that fills the masked positions in $\mathbf{y}^{p}$ using the assignment $\mathbf{z}$. The function $\mathsf{IsExtendable}(\cdot)$ is a standard capability provided by most grammar parsers (e.g., Earley (Earley, 1970) and LR (Aho and Johnson, 1974) parsers) and can efficiently check whether a complete prefix can be extended into a valid sentence.
Figure 2. The dLLM proposes a token for one masked position, which is subsequently verified through Lookahead-Based Verification. A proposed token is accepted if at least one lookahead complete prefix is extendable under the target grammar; otherwise, it is rejected.
3.2.2. Motivation and Feasibility.
dLLMs offer a distinct advantage over AR LLMs: they predict token distributions for all positions simultaneously. This property enables a natural lookahead-and-verify strategy. Given an incomplete prefix $\mathbf{y}^{p}$, we can perform lookahead by filling the masked positions in $\mathcal{M}(\mathbf{y}^{p})$ according to the model’s predicted distributions, thereby transforming $\mathbf{y}^{p}$ into a set of complete prefixes that reflect the model’s likely continuations. These complete prefixes can then be checked using a grammar parser via $\mathsf{IsExtendable}(\cdot)$, allowing us to verify whether the current output reliably remains extendable.
The effectiveness and efficiency of this strategy depend on how many complete prefixes need to be sampled. While sampling a larger number of complete prefixes increases the likelihood of acceptance, it may also introduce non-negligible overhead. To assess this trade-off and validate the feasibility of our approach, we conducted an empirical study.
We evaluated two widely used models, Dream-v0-Instruct-7B (Ye et al., 2025) and LLaDA-8B-Instruct (Nie et al., 2025), on the HumanEval-CPP benchmark (Zheng et al., 2023). For each test instance, we randomly select a prefix from the reference solution and prefill this complete prefix into the output. To simulate an intermediate inference state, we then randomly mask tokens within this complete prefix to obtain an incomplete prefix, which is guaranteed to be extendable. Specifically, each token is retained with a probability of 80%, and replaced with [MASK] otherwise. We then perform a single forward pass to obtain the predicted probability distributions at each masked position. To test the feasibility of our motivation, we sample $N$ distinct complete prefixes from these distributions and check whether any of them can be extended into a valid sentence under the target CFG. A higher acceptance rate indicates that our motivation is practically feasible in reliably determining the extendability of incomplete prefixes. As illustrated in Figure 3, we observe that sampling only a small number of complete prefixes is typically sufficient to uncover a valid complete prefix whenever the underlying incomplete prefix has the potential to be extended into one valid sentence. In particular, with a lookahead size of $N=10$, the acceptance rate reaches 98.1% for Dream-v0-Instruct-7B and 97.3% for LLaDA-8B-Instruct. These results indicate that the proposed lookahead-then-verify strategy is feasible in practice.
(a) Dream-v0-Instruct-7B
(b) LLaDA-8B-Instruct
Figure 3. Experimental results from our preliminary empirical study.
3.2.3. Verification Procedure.
Based on the feasibility of the proposed lookahead-then-verify strategy, we now describe how it is implemented in practice.
Let $P_{j}(\cdot)$ denote the predicted token distribution at each masked position $j\in\mathcal{M}(\mathbf{y}^{p})$. To perform lookahead, we sample $N$ distinct assignments $\{\mathbf{z}^{(1)},\ldots,\mathbf{z}^{(N)}\}$ according to these marginal distributions $P_{j}(\cdot)$. Formally, each assignment $\mathbf{z}$ is sampled from the joint distribution:
(4) $P(\mathbf{z})=\prod_{j\in\mathcal{M}(\mathbf{y}^{p})}P_{j}(z_{j}).$
Each sampled assignment $\mathbf{z}^{(n)}$ is then used to fill all masked positions in $\mathbf{y}^{p}$, resulting in a complete prefix:
(5) $\mathbf{s}^{(n)}=\Phi(\mathbf{y}^{p},\mathbf{z}^{(n)}),n\in\{1,\cdots,N\},$
which no longer contains any [MASK] tokens.
In the verification stage, we apply grammar parsers to each complete prefix to determine whether it is extendable in parallel. The proposed token is accepted if and only if at least one complete prefix satisfies:
(6) $\exists n\in\{1,\ldots,N\}\quad\text{s.t.}\quad\mathsf{IsExtendable}(\mathbf{s}^{(n)})=\texttt{true}.$
Such a complete prefix serves as a constructive witness that the updated output remains grammatically extendable, thereby naturally ensuring the reliability of the constraint. Conversely, if no valid complete prefix is found after $N$ lookahead attempts, the proposed token is rejected, and another proposed token is sampled from the model’s original predicted distribution.
3.3. Cache-Enhanced Recovery
In practice, the iterative proposal and verification process may occasionally enter a challenging state in which the current context makes it difficult for the model to produce an acceptable token. Concretely, the model may repeatedly propose tokens that fail the Lookahead-Based Verification, causing the decoding process to nearly stall. To address this issue, we introduce a recovery strategy that actively modifies the current context.
Our recovery strategy leverages a byproduct of the Lookahead-Based Verification process described in Section 3.2. Specifically, whenever a complete prefix $\mathbf{s}^{(n)}$ passes the verification, it constitutes a valid prefix that is both syntactically extendable and favored by the model. We record such prefixes by maintaining a global variable $\mathbf{y}_{\text{cache}}$, which is updated whenever a verification succeeds:
(7) $\mathbf{y}_{\text{cache}}\leftarrow\mathbf{s}^{(n)},\quad\text{if }\mathsf{IsExtendable}(\mathbf{s}^{(n)})=\texttt{true}.$
During inference, we monitor the number of consecutive failed proposal attempts. When this count exceeds a predefined proposal budget $\tau$, we infer that the model is trapped in a difficult context. To break this impasse, we replace the prefix of the current output with $\mathbf{y}_{\text{cache}}$:
(8) $\mathbf{y}_{1:r(\mathbf{y})}\leftarrow\mathbf{y}_{\text{cache}}.$
However, when $\mathbf{y}_{1:r(\mathbf{y})}$ contains no remaining [MASK] tokens (i.e., $\mathbf{y}_{1:r(\mathbf{y})}$ is identical to $\mathbf{y}_{\text{cache}}$), this replacement alone does not change the current context. To ensure that the context can be effectively modified, we generate one additional token following the current prefix. Many parsers used in existing constrained decoding approaches (Microsoft, 2025; Chen et al., 2025; Wagner and Graham, 1998) can efficiently compute the set of valid successor tokens given a complete prefix. All tokens in this set are guaranteed to preserve the syntactic validity of the prefix. Let $\mathcal{V}=\mathsf{NextTokens}(\mathbf{y}_{\text{cache}})$ denote the set of such admissible tokens. We define an adjusted probability distribution for the next position as follows:
(9) $\hat{P}_{r(\mathbf{y})+1}(t)\propto P_{r(\mathbf{y})+1}(t)\,\mathbb{I}[t\in\mathcal{V}].$
We then sample the next token from $\hat{P}_{r(\mathbf{y})+1}$ and continue inference with the standard proposal and verification process, thereby preserving both the non-autoregressive generation behavior and the constraint reliability of subsequent generation.
4. Experimental Setup
In this section, we introduce the experimental setup, including the benchmark and the corresponding context-free grammar, the compared baselines, the evaluation metrics, the models, and other implementation details. Based on this setup, we conduct a series of experiments designed to address the following Research Questions (RQs).
  • RQ1: How well does LAVE ensure the syntactic correctness of dLLM outputs compared to baseline approaches?
  • RQ2: How well does LAVE improve the functional correctness of dLLM outputs?
  • RQ3: How about the runtime overhead introduced by LAVE?
  • RQ4: How does each component of LAVE contribute to the overall performance?
  • RQ5: How do the hyperparameters affect the performance of LAVE?
4.1. Benchmark
To ensure a fair comparison, we follow existing studies (Mündler et al., 2025) in selecting benchmarks for evaluation. Specifically, we adopt three benchmarks that require the model to generate representative formal languages defined by context-free grammars.
  • Code Generation: HumanEval-CPP (Zheng et al., 2023) is constructed based on the HumanEval benchmark (Chen, 2021) and is designed to evaluate the code generation ability of LLMs in the C++ programming language. The benchmark contains 164 problems. For each problem, the dataset provides a natural language problem description and a set of test cases to verify the correctness of the output. We include this benchmark to demonstrate the performance of LAVE in code generation scenarios. In the following sections, we refer to this benchmark as CPP-Bench.
  • JSON Generation: JSON-Mode-Eval-Extended (Mündler et al., 2025) is constructed based on the JSON-Mode-Eval benchmark (NousResearch, 2024) to evaluate the information extraction capabilities of LLMs using the JSON format. The benchmark consists of 272 problems. Each problem comprises a natural language snippet, a JSON Schema (corresponding to a specific CFG), and a ground truth JSON response. We include this benchmark because it can effectively reflect the performance of LAVE in generating JSON, a formal language which is widely employed in information extraction and agent function calling tasks. In the following sections, we refer to this benchmark as JSON-Bench.
  • Chemical Expression Generation: SMILES-Eval (Mündler et al., 2025) is constructed to assess LLM capabilities in generating SMILES chemical molecule representations from natural language descriptions. The benchmark consists of 167 problems. Each problem includes a natural language molecule description and a corresponding SMILES string as the ground truth. We include this benchmark to highlight the applicability of LAVE in scientific domains. In the following sections, we refer to this benchmark as SMILES-Bench.
4.2. Grammar
For CFG-constrained decoding, a concrete grammar specification is required. For CPP-Bench and SMILES-Bench, we construct grammars based on the standard syntax of C++ and SMILES, respectively. For JSON-Bench, where each problem is associated with a distinct JSON schema, we construct a separate grammar for each problem.
4.3. Baselines
We compare LAVE against three baseline approaches. ❶ First, we include the constrained decoding approach proposed by Mündler et al. (Mündler et al., 2025), which we refer to as IG-CD (Intersection-Guided Constrained Decoding). ❷ Second, we include the unconstrained generation setting, denoted as NO-CD. ❸ In addition, we design a straightforward constrained decoding baseline that enforces a strictly left-to-right generation order for dLLMs, thereby allowing the direct application of constrained decoding techniques (Microsoft, 2025) originally developed for AR LLMs. We refer to this baseline as FS-CD (Forced-Sequential Constrained Decoding).
4.4. Metrics
We evaluate all approaches using three metrics: syntactic@k, functional@k, and average inference time.
syntactic@k measures the proportion of problems for which at least one of the $k$ generated outputs satisfies the syntactic constraints specified by the given context-free grammar.
functional@k measures the proportion of problems for which at least one of the $k$ generated outputs exhibits correct functional behavior. For CPP-Bench, functional correctness is determined by executing the provided test cases. For the other benchmarks, it is assessed by exact match against the ground-truth outputs. In general, syntactic correctness is a prerequisite for functional correctness.
Average Inference Time records the average decoding time per generation, providing a measure of the overhead introduced by different constrained decoding approaches.
4.5. Models
We evaluate our approach using four widely used dLLMs:
  • LLaDA-8B-Instruct (Nie et al., 2025) is released by researchers from Renmin University of China and Ant Group. It is an early open-source diffusion large language model with 8 billion parameters and trained entirely from scratch. In the following, we refer to this model as LLaDA-8B.
  • LLaDA-1.5 (Zhu et al., 2025) is also released by researchers from Renmin University of China and Ant Group. It extends LLaDA-8B-Instruct by incorporating Variance Reduced Preference Optimization (VRPO) during training. Compared with LLaDA-8B-Instruct, it achieves improved performance across a wide range of tasks, including code-related tasks.
  • Dream-v0-Instruct-7B (Ye et al., 2025) is released by researchers from The University of Hong Kong and Huawei. It is initialized from an autoregressive language model and demonstrates strong overall capabilities. In the following, we refer to this model as Dream-7B.
  • DiffuCoder-7B-cpGRPO (Gong et al., 2025) is released by researchers from Apple and The University of Hong Kong. It is post-trained with coupled-GRPO on 21K code examples, leading to substantial improvements in code generation performance. In the following, we refer to this model as DiffuC-7B.
4.6. Implementation Details
Following prior studies (Yang et al., 2025a; Mündler et al., 2025), we apply a simple and commonly used optimization for all approaches: once a dLLM generates an [EOS] token, all subsequent token positions are set to [EOS] to accelerate inference. For all experiments, we adopt the widely used semi-autoregressive strategy (Nie et al., 2025; Arriola et al., 2025; Yang et al., 2025b; Wang et al., 2025a, b) and set the block size to 32, as is common in prior studies (Zhu et al., 2025; Nguyen-Tri et al., 2025; Hu et al., 2025). What’s more, for each experiment, we conduct at least four independent runs and report the average values of the evaluation metrics. For both FS-CD and LAVE, we employ the Earley algorithm (Earley, 1970) to implement the required grammar parser, which allows us to handle arbitrary context-free grammars. All experiments are conducted on a server equipped with eight NVIDIA A100 GPUs with 40 GB memory each, and two Intel Xeon Gold 6348 CPUs, with 28 cores per socket and 56 physical cores in total.
Regarding the hyperparameters of the dLLMs, we set the max generation length $L$ to 256. The number of denoising steps $T$ is set to 128, which provides a balanced trade-off between model effectiveness and efficiency (Li et al., 2025a; Nie et al., 2025). The temperature is fixed to 0.2 for all experiments. Regarding the hyperparameters specific to LAVE, we set lookahead size $N$ defined in Section 3.2 to 10 and the attempt budget $\tau$ defined in Section 3.3 to 5.
Additional implementation details, including the grammar used for each task and the specific implementations of compared methods, are provided in the Supplementary Material.
5. Experimental Results
Figure 4. Outputs generated by NO-CD, FS-CD, IG-CD, and LAVE on Dream-7B for a problem from CPP-Bench, respectively.
5.1. RQ1: How Well Does LAVE Ensure the Syntactic Correctness of dLLM Outputs?
Motivation. The primary objective of LAVE is to ensure that the outputs generated by dLLMs reliably conform to the syntax defined by a CFG. To assess this capability, we evaluate the syntactic correctness of the outputs generated by our approach compared to baselines in this RQ.
Setting. We apply the baselines and LAVE to the four models described in Section 4.5 and evaluate their performance on the three benchmarks introduced in Section 4.1. We use syntactic@k as the evaluation metric with $k\in\{1,3,5,10\}$, as defined in Section 4.4.
Table 1. syntactic@k (%) on JSON-Bench, CPP-Bench, and SMILES-Bench, where $k\in\{1,3,5,10\}$. The best result for each model, benchmark, and $k$ is highlighted in bold, while the second-best result is underlined.
Method JSON-Bench CPP-Bench SMILES-Bench Average
k=1 k=3 k=5 k=10 k=1 k=3 k=5 k=10 k=1 k=3 k=5 k=10 k=1 k=3 k=5 k=10
LLaDA-8B
NO-CD 85.4 88.9 90.1 96.4 74.2 90.5 94.1 95.6 66.0 74.1 75.9 78.9 75.2 84.5 86.7 90.3
FS-CD 94.9 95.7 96.4 96.4 66.3 66.7 66.7 66.7 75.7 75.7 75.9 76.1 79.0 79.4 79.7 79.7
IG-CD 94.8 95.1 95.7 96.7 86.1 92.6 96.7 98.1 94.3 98.1 98.7 99.2 91.7 95.2 97.0 98.0
LAVE 99.5 99.6 99.6 99.6 90.9 98.7 99.3 99.4 98.7 99.4 100.0 100.0 96.4 99.2 99.6 99.7
LLaDA-1.5
NO-CD 86.0 87.1 87.7 87.9 78.2 89.6 92.0 94.8 62.3 74.1 77.7 83.1 75.5 83.6 85.8 88.6
FS-CD 96.0 97.2 97.7 98.1 73.8 74.0 74.2 74.2 78.3 78.5 78.5 78.5 82.7 83.2 83.5 83.6
IG-CD 94.0 95.1 95.4 96.2 89.5 96.1 97.3 98.8 93.7 96.7 97.0 97.6 92.4 96.0 96.6 97.5
LAVE 99.7 99.7 99.8 99.8 91.3 97.6 98.8 98.8 96.6 99.3 100.0 100.0 95.9 98.9 99.5 99.5
Dream-7B
NO-CD 86.8 88.9 89.3 89.3 76.2 76.8 79.2 81.1 74.0 75.6 76.2 76.8 79.0 80.4 81.6 82.4
FS-CD 96.8 97.9 98.2 98.2 83.1 84.3 84.4 84.4 93.1 93.8 93.8 93.8 91.0 92.0 92.1 92.1
IG-CD 94.9 95.3 95.5 95.6 87.6 90.2 90.4 90.9 95.1 96.2 96.7 97.4 92.5 93.9 94.2 94.6
LAVE 99.0 99.3 99.6 99.7 96.0 98.1 99.4 99.4 99.4 99.7 100.0 100.0 98.1 99.0 99.7 99.7
DiffuC-7B
NO-CD 89.9 90.2 90.4 90.4 79.9 80.2 80.5 81.7 70.1 71.1 71.3 71.3 80.0 80.5 80.7 81.1
FS-CD 96.9 97.3 98.1 98.1 93.8 94.0 94.0 94.3 96.0 96.0 96.3 96.3 95.6 95.8 96.1 96.2
IG-CD 92.1 92.5 92.7 92.7 93.8 94.5 95.1 95.3 96.4 96.6 96.6 96.7 94.1 94.5 94.8 94.9
LAVE 99.6 100.0 100.0 100.0 97.0 99.1 99.6 100.0 98.8 99.1 99.3 99.4 98.6 99.4 99.6 99.8
Results. The syntactic@k of different approaches is reported in Table 1.
LAVE demonstrates superior performance in ensuring that dLLM outputs conform to CFG constraints. Compared to the unconstrained baseline, LAVE substantially improves syntactic correctness, reaching close to 100 percent in most settings. For example, on the syntactically simpler SMILES-Bench with LLaDA-8B, LAVE raises syntactic@1 from 66.0% to 98.7%, syntactic@3 from 74.1% to 99.4%, syntactic@5 from 75.9% to 100.0%, and syntactic@10 from 78.9% to 100.0%. On the more syntactically complex CPP-Bench, our approach also achieves strong performance. With DiffuC-7B, LAVE increases syntactic@1 from 79.9% to 97.0%, syntactic@3 from 80.2% to 99.1%, syntactic@5 from 80.5% to 99.6%, and syntactic@10 from 81.7% to 100.0%. These results demonstrate that LAVE greatly enhances the reliability of dLLM outputs when generating formal languages such as C++, JSON, and SMILES.
LAVE substantially outperforms existing constrained decoding approaches across most benchmarks and models. For example, on JSON-Bench with LLaDA-8B, LAVE improves syntactic@1 over the best baseline FS-CD by 4.6%. On CPP-Bench with Dream-7B, LAVE outperforms the best baseline IG-CD by 8.4% in terms of syntactic@1. In addition, we observe that LAVE benefits more from increasing the $k$ than FS-CD. For instance, on CPP-Bench with LLaDA-1.5, as $k$ increases from 1 to 10, LAVE improves syntactic@k from 91.3% to 98.8%, whereas FS-CD shows only a marginal increase from 73.8% to 74.2%. We attribute this behavior to the fact that our approach preserves the non-autoregressive generation nature of dLLMs, which provides greater flexibility and diversity during inference.
We further conduct a case study to analyze why LAVE outperforms existing approaches. Figure 4 presents a representative example in which only LAVE produces a syntactically valid output. In this example, the inference process of IG-CD stalls with a single remaining [MASK] token. Theoretically, completing the output at this point requires multiple tokens (such as a variable name, a closing parenthesis, and a semicolon). However, since only one masked position remains, it becomes impossible for the model to generate a grammatically correct completion. We attribute this failure to the unreliable constraints imposed by IG-CD, which allow the model to enter a state where valid continuation is no longer possible. In contrast, FS-CD suffers from severe repetition. As shown in Figure 4, the model repeatedly generates the same if statement until the output reaches the max generation length, resulting in a forced termination. We attribute this issue to FS-CD’s strict enforcement of left-to-right generation, which disrupts the intrinsic non-autoregressive generation nature of dLLMs and significantly degrades the quality of the predicted token distributions.
Answer to RQ1: LAVE effectively enforces CFG constraints on generated outputs and substantially outperforms all baselines across most benchmarks and models. Compared to existing approaches, LAVE reliably ensures the validity of every newly generated token throughout the inference process.
5.2. RQ2: How Well Does LAVE Improve the Functional Correctness of dLLM Outputs?
Motivation. Although the primary objective of CFG-constrained decoding is to guarantee the syntactic correctness of generated outputs, we are also interested in whether LAVE can improve the functional correctness of generated outputs. In this RQ, we evaluate the impact of LAVE on the functional correctness of dLLM outputs.
Setting. We apply the baselines and LAVE to the four dLLMs described in Section 4.5. We then measure the performance of different approaches on JSON-Bench and CPP-Bench. We use functional@k as the evaluation metric with $k\in\{1,3,5,10\}$, as defined in Section 4.4. We exclude SMILES-Bench because all models yielded a 0% functional correctness rate on this benchmark, making meaningful comparison infeasible.
Table 2. functional@k (%) on JSON-Bench and CPP-Bench, where $k\in\{1,3,5,10\}$. The best result for each model, benchmark, and $k$ is highlighted in bold. Relative improvements are computed with respect to NO-CD.
Method JSON-Bench CPP-Bench
k=1 k=3 k=5 k=10 k=1 k=3 k=5 k=10
LLaDA-8B
NO-CD 46.9 (+0.0) 48.8 (+0.0) 49.2 (+0.0) 49.6 (+0.0) 16.3 (+0.0) 21.8 (+0.0) 23.1 (+0.0) 24.9 (+0.0)
FS-CD 52.9 (+6.0) 53.1 (+4.3) 53.3 (+4.1) 53.5 (+3.9) 15.7 (-0.6) 16.0 (-5.8) 16.0 (-7.1) 16.2 (-8.7)
IG-CD 52.1 (+5.2) 52.9 (+4.1) 53.3 (+4.1) 53.7 (+4.1) 17.8 (+1.5) 22.5 (+0.7) 23.7 (+0.6) 25.6 (+0.7)
LAVE 55.0 (+8.1) 56.1 (+7.3) 56.3 (+7.1) 56.4 (+6.8) 19.5 (+3.2) 23.8 (+2.0) 25.9 (+2.8) 26.7 (+1.8)
LLaDA-1.5
NO-CD 45.5 (+0.0) 46.2 (+0.0) 46.5 (+0.0) 46.7 (+0.0) 17.2 (+0.0) 20.1 (+0.0) 22.0 (+0.0) 25.9 (+0.0)
FS-CD 53.1 (+7.6) 53.3 (+7.1) 53.6 (+7.1) 53.6 (+6.9) 15.9 (-1.3) 16.1 (-4.0) 16.5 (-5.5) 16.5 (-9.4)
IG-CD 50.6 (+5.1) 51.8 (+5.6) 52.2 (+5.8) 52.5 (+5.8) 18.3 (+1.1) 21.3 (+1.2) 23.1 (+1.1) 27.1 (+1.2)
LAVE 54.1 (+8.6) 54.5 (+8.3) 54.7 (+8.2) 54.7 (+8.0) 19.1 (+1.9) 23.0 (+2.9) 24.9 (+2.9) 28.8 (+2.9)
Dream-7B
NO-CD 49.8 (+0.0) 50.4 (+0.0) 50.7 (+0.0) 50.7 (+0.0) 25.6 (+0.0) 26.2 (+0.0) 26.8 (+0.0) 28.1 (+0.0)
FS-CD 52.9 (+3.1) 53.1 (+2.7) 53.2 (+2.5) 53.2 (+2.5) 26.5 (+0.9) 26.8 (+0.6) 27.1 (+0.3) 27.5 (-0.6)
IG-CD 51.5 (+1.7) 53.3 (+2.9) 53.5 (+2.8) 53.9 (+3.2) 29.1 (+3.5) 30.5 (+4.3) 30.8 (+4.0) 31.1 (+3.0)
LAVE 54.4 (+4.6) 54.7 (+4.3) 55.1 (+4.4) 55.5 (+4.8) 33.5 (+7.9) 38.2 (+12.0) 39.5 (+12.7) 41.4 (+13.3)
DiffuC-7B
NO-CD 52.0 (+0.0) 52.9 (+0.0) 53.3 (+0.0) 53.7 (+0.0) 37.2 (+0.0) 37.4 (+0.0) 37.8 (+0.0) 37.8 (+0.0)
FS-CD 53.7 (+1.7) 54.1 (+1.2) 54.1 (+0.8) 54.1 (+0.4) 35.2 (-2.0) 35.5 (-1.9) 35.8 (-2.0) 35.9 (-1.9)
IG-CD 53.0 (+1.0) 54.7 (+1.8) 55.3 (+2.0) 55.3 (+1.6) 38.1 (+0.9) 38.3 (+0.9) 38.3 (+0.5) 38.5 (+0.7)
LAVE 55.2 (+3.2) 55.5 (+2.6) 56.6 (+3.3) 56.7 (+3.0) 38.1 (+0.9) 38.4 (+1.0) 38.5 (+0.7) 38.5 (+0.7)
Results. The functional@k of different approaches is reported in Table 2.
LAVE consistently improves the functional correctness of dLLMs across four models and two benchmarks. Compared to the unconstrained baseline (NO-CD), LAVE achieves higher functional@k in all settings. In contrast, other constrained decoding baselines sometimes degrade model performance. For example, on CPP-Bench with LLaDA-8B, FS-CD reduces functional@1 by 0.6%, whereas LAVE increases it by 3.2%.
LAVE outperforms other CFG constrained decoding approaches in improving functional correctness. The gains achieved by LAVE are substantially larger than those of the baseline methods. For example, on Dream-7B with CPP-Bench, LAVE improves functional@1, functional@3, functional@5, and functional@10 from 25.6%, 26.2%, 26.8%, and 28.1% under the unconstrained setting to 33.5%, 38.2%, 39.5%, and 41.4%, respectively. In contrast, the best baseline reaches only 29.1%, 30.5%, 30.8%, and 31.1%. This trend remains consistent across other experimental settings.
LAVE yields larger functional correctness gains on tasks with limited reasoning requirements. JSON-Bench focuses on structured information extraction and typically does not require complex logical reasoning, whereas CPP-Bench involves code generation and is more dependent on reasoning ability. We observe that, compared to NO-CD, LAVE achieves larger improvements in functional@k on JSON-Bench than on CPP-Bench. For example, with LLaDA-8B, LAVE improves functional correctness by 8.1% on JSON-Bench, whereas the improvement is 3.2% on CPP-Bench. This behavior is expected because, in tasks with limited reasoning requirements, functional correctness depends more strongly on syntactic validity, allowing the syntactic improvements provided by LAVE to translate more effectively into functional gains.
Answer to RQ2: Compared to baselines, LAVE improves the functional correctness of dLLMs more consistently and significantly. Furthermore, this improvement is particularly evident in tasks with lower logical reasoning requirements.
5.3. RQ3: How about the Runtime Overhead Introduced by LAVE?
Motivation. Beyond the quality of generated outputs, runtime efficiency is also an important factor that influences the practical adoption of constrained decoding approaches. Therefore, in this RQ, we examine the runtime overhead introduced by LAVE.
Table 3. Average inference time (s) on JSON-Bench, CPP-Bench, and SMILES-Bench. The best result for each model and benchmark is highlighted in bold, while the second-best result is underlined. The last row reports the relative inference time normalized by NO-CD.
Model JSON-Bench CPP-Bench SMILES-Bench
NO-CD FS-CD IG-CD LAVE NO-CD FS-CD IG-CD LAVE NO-CD FS-CD IG-CD LAVE
LLaDA-8B 9.48 13.29 10.34 8.22 8.55 16.68 9.66 9.30 8.49 7.68 8.76 3.02
LLaDA-1.5 9.71 14.74 10.82 9.91 8.56 11.88 12.17 12.02 8.77 2.16 8.21 2.97
Dream-7B 8.62 15.74 8.25 9.73 7.46 7.17 7.95 7.26 7.07 8.99 19.64 2.24
DiffuC-7B 8.15 15.39 9.78 9.22 7.45 14.78 8.47 12.27 6.86 11.15 6.74 4.41
Average 8.99 14.79 9.80 9.27 8.01 12.63 9.56 10.21 7.80 7.50 10.84 3.16
Ratio $\times$1.65 $\times$1.09 $\times$1.03 $\times$1.58 $\times$1.19 $\times$1.27 $\times$0.96 $\times$1.39 $\times$0.41
Setting. We apply the baselines and LAVE to the four models described in Section 4.5 and evaluate them on the three benchmarks introduced in Section 4.1. We use average inference time as the metric to measure the runtime overhead of different approaches.
Results. The average inference times of different approaches are reported in Table 3. Lower values indicate better efficiency.
LAVE introduces negligible runtime overhead across most benchmarks and models. For example, on LLaDA-1.5 evaluated on JSON-Bench, applying LAVE increases the average inference time from 9.71 seconds to 9.91 seconds, corresponding to an increase of only 2%. For LLaDA-8B on CPP-Bench, where the grammar of C++ is considerably more complex than that of JSON or SMILES, the inference time increases modestly from 8.55 seconds to 9.30 seconds, corresponding to a 9% increase. Interestingly, we observe that in several scenarios, the average inference time of LAVE is even lower than that of the unconstrained baseline. For example, with DiffuC-7B on SMILES-Bench, the average inference time for LAVE is 4.41 seconds, compared to 6.86 seconds for the unconstrained setting. We attribute this phenomenon to the observation that, without constraints, the model often generates syntactically invalid outputs (such as extraneous natural language explanations), which results in the generation of a larger number of tokens and consequently prolongs the inference process.
Answer to RQ3: LAVE introduces negligible runtime overhead in most settings, allowing it to improve the reliability of structured generation in dLLMs with minimal additional cost.
5.4. RQ4: How Does Each Component of LAVE Contribute to the Overall Performance?
Table 4. Ablation study of LAVE. Verification and Recovery denote Lookahead-Based Verification and Cache-Enhanced Recovery, respectively. Syn and Func represent syntactic@1 and functional@1, respectively, while Time denotes the average inference time.
Model Validation Recovery JSON-Bench CPP-Bench
Syn. (%) Func. (%) Time (s) Syn. (%) Func. (%) Time (s)
LLaDA-8B 85.4 (+0.0) 46.9 (+0.0) 9.48 74.2 (+0.0) 16.3 (+0.0) 8.55
95.8 (+10.4) 53.7 (+6.8) 8.09 86.6 (+12.4) 18.9 (+2.6) 8.71
99.5 (+14.1) 55.0 (+8.1) 8.22 90.9 (+16.7) 19.5 (+3.2) 9.30
Dream-7B 86.8 (+0.0) 49.8 (+0.0) 8.62 76.2 (+0.0) 25.6 (+0.0) 7.46
94.8 (+8.0) 51.8 (+2.0) 9.17 92.7 (+16.5) 32.3 (+6.7) 7.69
99.0 (+12.2) 54.4 (+4.6) 9.73 96.0 (+19.8) 33.5 (+6.9) 7.26
Motivation. LAVE consists of two components, namely Lookahead-Based Verification and Cache-Enhanced Recovery. In this RQ, we evaluate the contribution of each component through ablation experiments.
Setting. We conduct experiments on LLaDA-8B and Dream-7B using JSON-Bench and CPP-Bench. All experiments are evaluated using the three metrics defined in Section 4.4. Since Cache-Enhanced Recovery depends on Lookahead-Based Verification, it is not possible to remove Lookahead-Based Verification while retaining Cache-Enhanced Recovery. Therefore, we consider three experimental settings: ❶ retaining both Lookahead-Based Verification and Cache-Enhanced Recovery, corresponding to LAVE; ❷ retaining Lookahead-Based Verification while removing Cache-Enhanced Recovery, where the model switches to unconstrained generation once the triggering condition for Cache-Enhanced Recovery is met; ❸ removing both Lookahead-Based Verification and Cache-Enhanced Recovery, which corresponds to NO-CD.
Results. The results of the ablation experiments are reported in Table 4.
Lookahead-Based Verification alone significantly improves both syntactic and functional correctness. For example, on LLaDA-8B evaluated on JSON-Bench, applying Lookahead-Based Verification increases the syntactic@1 by 10.4% and the functional@1 by 6.8%. On Dream-7B evaluated on CPP-Bench, applying Lookahead-Based Verification increases the syntactic@1 by 16.5% and the functional@1 by 6.7%.
Cache-Enhanced Recovery further improves the performance of LAVE without introducing additional runtime overhead. Across all evaluated settings, Cache-Enhanced Recovery consistently improves both syntactic and functional correctness compared to using Lookahead-Based Verification alone, while having a negligible impact on average inference time. For example, on Dream-7B evaluated on JSON-Bench, it increases the syntactic@1 to 99.0% and the functional@1 to 54.4%, with an added runtime cost of less than 0.6 seconds.
Answer to RQ4: Both Lookahead-Based Verification and Cache-Enhanced Recovery contribute effectively to the performance of LAVE by improving the syntactic and functional correctness of dLLM outputs.
5.5. RQ5: How Do the Hyperparameters Affect the Performance of LAVE?
(a) lookahead size NN
(b) proposal budget $\tau$
Figure 5. The performance of LAVE with different values of the lookahead size $N$ and proposal budget $\tau$. We use the gray dashed line to represent the employed hyperparameters.
Table 5. Performance of LAVE with varying values of the hyperparameter $T$. The gray background indicates the hyperparameter setting used in our experiments.
Method syntactic@1 (%) functional@1 (%) Average Inference Time (s)
T=16T{=}16 T=32T{=}32 T=64T{=}64 T=128T{=}128 T=256T{=}256 T=16T{=}16 T=32T{=}32 T=64T{=}64 T=128T{=}128 T=256T{=}256 T=16T{=}16 T=32T{=}32 T=64T{=}64 T=128T{=}128 T=256T{=}256
NO-CD 7.8 21.0 63.6 86.8 97.4 1.1 7.4 40.4 49.8 54.8 1.98 2.92 4.56 8.62 15.88
LAVE 76.3 91.7 97.2 99.0 100.0 20.4 32.2 46.6 54.4 54.8 3.21 3.76 5.48 9.73 16.02
Motivation. LAVE involves several hyperparameters that may affect its performance. In this RQ, we conduct a sensitivity analysis to examine their impact. Specifically, we consider two categories of hyperparameters. ➀ The first category consists of hyperparameters introduced by LAVE, including the lookahead size $N$, defined in Section 3.2, and the attempt budget $\tau$, defined in Section 3.3. ➁ The second category includes key hyperparameters inherent to dLLMs, for which we select the number of denoising steps $T$.
Setting. We conduct experiments on Dream-7B using JSON-Bench. All experiments are evaluated using the three metrics defined in Section 4.4. For the lookahead size $N$, we consider values in $\{3,5,10,20,30,40\}$. For the attempt budget $\tau$, we consider values in $\{1,3,5,10,20,30\}$. For the number of denoising steps $T$, we consider values in $\{16,32,64,128,256\}$.
Results.
The results for the hyperparameters introduced by LAVE, namely the lookahead size $N$ and the attempt budget $\tau$, are shown in Figure 5.Effect of the lookahead size $N$. We observe that increasing lookahead size $N$ generally leads to improvements in both syntactic@1 and functional@1. This trend can be attributed to the fact that a larger lookahead size $N$ increases the likelihood of identifying at least one valid complete prefix during Lookahead-Based Verification. Interestingly, the average inference time first decreases and then increases as $N$ grows. This behavior reflects two competing effects. On the one hand, a larger $N$ increases the number of complete prefixes evaluated during each validation step, which increases inference time. On the other hand, a larger $N$ increases the probability that a valid proposed token is accepted, thereby reducing the number of re-proposal attempts and shortening the overall inference time. ❷ Effect of the proposal budget $\tau$. As $\tau$ increases, all three metrics exhibit an upward trend. A larger attempt budget allows the model more opportunities to propose high-confidence tokens, which improves both syntactic and functional correctness. At the same time, increasing $\tau$ leads to more re-proposal, which increases inference time. ❸ Overall robustness to hyperparameter variations. Overall, LAVE demonstrates strong robustness to adjustments in the newly introduced hyperparameters. Although varying lookahead size $N$ and $\tau$ leads to observable performance trends, the magnitude of these changes remains limited. For example, when $N$ increases from 3 to 40, the average inference time remains stable at around 10 seconds. Similarly, when $\tau$ increases from 3 to 30, the syntactic@1 consistently remains above 97%.
The results for the hyperparameters inherent to dLLMs, namely the denoising steps $T$, are shown in Table 5.LAVE consistently improves syntactic and functional correctness across different values of $T$. For instance, when $T$ is set to 32, which in our setting corresponds to generating at least eight tokens per forward pass, LAVE increases the syntactic@1 from 21.0% to 91.7% and the functional@1 from 7.4% to 32.2%. Although the runtime overhead increases slightly when the number of denoising steps is small (e.g., rising from 2.92s to 3.76s at $T=32$), we consider this cost acceptable given the substantial gains in correctness. ❷ The syntactic and functional correctness of dLLMs with LAVE increases as $T$ increases. For instance, as $T$ increases from 16 to 128, the syntactic@1 improves from 76.3% to 99.0%, and the functional@1 rises from 20.4% to 54.4%. We attribute this to the characteristics of dLLMs, where a higher number of denoising steps leads to better effectiveness but lower efficiency. The unconstrained baseline exhibits a similar pattern, further supporting this explanation.
Answer to RQ5: LAVE exhibits strong robustness to variations in key hyperparameters and consistently improves the quality of generated outputs across different settings.
6. Discussion
6.1. Understanding the Remaining Syntax Errors
Although our approach guarantees reliable grammar constraint by ensuring that no newly generated token renders the intermediate output impossible to complete into a valid sentence, occasional syntax errors still occur (Section 5.1). We identify the root cause of these remaining errors as the model’s tendency to generate repetitive patterns, a phenomenon widely observed in previous studies (Fu et al., 2021; Holtzman et al., 2019; Li et al., 2023). In such cases, the dLLM repetitively generates specific content (e.g., whitespace or newline characters) until it reaches the max generation length. To confirm this, we examined all invalid outputs from our experiments and found that the number of generated tokens exactly matched the max generation length limit in every case. Fortunately, the majority of approaches (Dong et al., 2025; Li et al., 2023; Xu et al., 2022) designed to resolve such repetition problems are orthogonal to our approach; therefore, they can be effectively combined with LAVE to further improve the reliability of dLLM generation.
6.2. Threats to Validity
We identify four primary threats to the validity of our study:
Generalizability of the Experimental Results. A potential concern is whether our findings hold across diverse tasks and models. To mitigate this threat, we carefully selected a broad range of benchmarks and models for our evaluation. Specifically, we adopted three representative benchmarks covering code generation, information extraction, and scientific applications. These domains reflect typical real-world scenarios where structured outputs are essential, aligning with setups in prior research (Mündler et al., 2025). Furthermore, we evaluated our approach on four distinct and widely used dLLMs, ensuring that our conclusions are not artifacts of a specific model architecture or training data distribution. Regarding evaluation, we employed three complementary metrics: syntactic@k, functional@k, and average inference time. Together, these metrics provide a holistic assessment, capturing not only the generation quality but also the practical computational cost of our approach.
Reliability-Oriented Constraint Design. Due to the non-autoregressive generation nature of dLLMs, it is inherently difficult to guarantee that all accepted tokens are valid while all rejected tokens are invalid. As a result, our approach may occasionally reject proposed tokens that could still lead to valid completions. Nevertheless, we argue that the primary goal of constrained decoding is to improve the reliability of generated outputs. From this perspective, ensuring that all accepted tokens preserve grammatical validity is more important than accepting every potentially valid token. Empirically, Section 3.2.2 suggests that a small lookahead budget is usually sufficient to uncover an extendable candidate when one exists, reducing the likelihood of rejecting valid tokens. The results in Section 5.1 and Section 5.2 further show that our method consistently improves syntactic correctness and also increases functional correctness, indicating that the practical impact of rejecting some valid tokens is negligible.
Data Leakage. The extensive pre-training corpora used for dLLMs, typically sourced from open-source communities, may inadvertently contain samples from our evaluation benchmarks. While this is a common challenge in LLM research, we argue that it does not compromise the fairness of our comparative analysis. In our experiments, we apply both the baselines and our proposed approach to the identical set of dLLMs. Consequently, the relative improvements demonstrated by our approach over the baselines remain valid and credible indicators of its effectiveness. To further alleviate concerns regarding data leakage, we plan to incorporate more recent benchmarks in future work.
Impact of Hyperparameters. Our approach introduces two hyperparameters, which could potentially limit utility if the approach requires extensive tuning. To address this, we initially selected reasonable values based on intuition and observed consistent performance improvements across multiple benchmarks and models. To rigorously examine the sensitivity of these parameters, we conducted a comprehensive sensitivity analysis by varying their values over a wide range. The empirical results indicate that our approach demonstrates strong robustness to variations in hyperparameter settings, maintaining high performance without the need for precise tuning.
7. Conclusion
In this paper, we present LAVE, a new constrained decoding approach tailored to emerging diffusion LLMs. LAVE exploits the unique ability of these models to predict token distributions for all positions in parallel during a single forward pass. Our approach addresses the limitations of existing constrained decoding methods by ensuring reliable grammar constraints that reliably preserve the potential for intermediate outputs to be extended into valid sentences. Experimental results on code generation, JSON generation, and chemical expression generation demonstrate that LAVE can substantially improve the reliability of dLLMs in producing CFG-compliant outputs.
Data Availability
We will release the source code upon acceptance.